# Cross-domain Multi-modal Few-shot Object Detection via Rich Text

### v1

@@ 4.5 Ablation Study
We first verify that the length of the rich text sentences have a distinct impact on the model’s performance. To test ‘length’, we conduct experiments on the UODD dataset using 4 cases of text: (i) the raw category name, (ii) our manually-built rich text referencing Wikipedia, (iii) an extended version of (ii) by adding image context, (iv) LLM-generated rich text. Cases (ii) to (iv) all result in long sentences that have rich semantics, and thus are generally beneficial for the model performance. See Tab. 3 for our examples and quantitative results. However, not all extensions on the sentence length are effective. For example, in case (iii), our extended rich text for sea cucumbers has a slightly negative impact over the non-extended version, decreasing performance from 17.4 to 16.9. This may be because the extra image context with an distinct name of another category in UODD (i.e. sea urchin) has confused the model.
In addition, we analyze the effectiveness of our proposed multi-modal aggregated feature module and the rich semantic rectify module. As seen in Tab. 4, comparing to the single-modal baseline, our method can achieve significant improvement through using rich text multi-modal information. Our rich text rectify module also leads to an additional improvement.
Table 3: The effect of different language modalities.
Type of text
Example
UODD
None
e.g. None
9.2
Category name
e.g. Sea cucumbers
9.9
Rich text
e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles
17.4
Extended rich text
e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles; usually seen together with sea urchins.
16.9
LLM-Rich text
e.g. Sea cucumbers are marine invertebrates known for their elongated, leathery bodies, which are typically covered in spines or tentacles and lack a distinct head or tail
17.5
Table 4: The effect of different modules.
@@ 4.6 Visualization
In Fig. 5, we present representative qualitative detection results on the three CD-FSOD benchmarks. In these examples, the objective is to detect araneae (ArTaxOr), sea urchins and sea cucumbers (UODD), and windmills (DIOR). The inference uses a confidence threshold setting of 0.3 for all methods. Compared to baseline methods, we show qualitative improvements in a more accurate bounding box (first row), fewer false positives (second row), and improved detection confidence (third row). See the Appendix for additional qualitative results.
@@ 4.7 Discussion
We observe that it is desirable when the rich text description of the target objects are available, but the specific context of the rich text plays a critical role. Consequently, in practice, the user must carefully evaluate the rich text and substantially tune it for greater improvement. This might become a burden in some cases when (i) the user has a rich text, but it contains many descriptions that are irrelevant to the visual information, or (ii) the user does not have rich text and thus need to build one from scratch. For the first case, a general suggestion for practical usage is to keep the information most related to the appearance of the target objects. For the second case, a user could extend and improve upon our LLM-generated text approach. In future work, for example, we plan to train an discriminator to select the best rich text from multiple candidates.

### v2

@@ 4.1 Implementation Details
4.2 Datasets
Following the standard evaluation on the CD-FSOD benchmark provided by Xiong et al. [48], we use ArTaxOr [9], DIOR [28] and UODD [20] as our target domain few-shot datasets. ArTaxOr contains 7 arthropods categories, DIOR consists of 20 satellite images categories, and UODD has 3 underwater creature categories.
Type of text Example UODD
None e.g. None 9.2
Category name e.g. Sea cucumbers 9.9
Rich text e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles 17.4
Extended rich text e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles; usually seen together with sea urchins. 16.9
LLM-Rich text e.g. Sea cucumbers are marine invertebrates known for their elongated, leathery bodies, which are typically covered in spines or tentacles and lack a distinct head or tail 17.5
Table 3: The effect of different language modalities.
To measure the generalization of our method, we evaluate on the general FSOD benchmark of PASCAL VOC. Following the pioneering evaluation pipeline for FSOD [47], 15 of the categories of PASCAL VOC are selected as base classes for pre-training, and the remaining 5 categories are used as the novel classes for few-shot fine-tuning.
4.3 Results on CD-FSOD Datasets
As introduced in  Sec. 3.2, we build rich text manually or by using an LLM [44]. As indicated in  Tab. 1, either way could generate significant improvement on the out-of-domain benchmarks. It is observed that LLM-generated rich text brings better performance on these three benchmarks, because the generated text contains more comprehensive sentences. However, this is not a universal situation, because sometimes the LLM failed to generate distinct descriptions for the target categories when their names are abstract and technical.
4.4 Results on Standard FSOD Datasets
As an auxiliary analysis and to further evaluate the generalization of our method, we test our model on a common FSOD benchmark, i.e. few-shot PASCAL VOC, in  Tab. 2. Following TFA [47], as described in  Sec. 4.2, there are three fixed groups of base-novel category splits, namely split 1, 2, and 3 to avoid sampling bias. For each data split, the $n$-shot ($n\in\{1,2,3,5,10\}$) sampling for the novel categories are also pre-defined. The model is evaluated on these three respective dataset splits. In addition, as described in  Sec. 3.1, the base categories are also trained during fine-tuning to avoid forgetting, but have different sampling strategies: (i) in works such as TFA [47], the base categories also follow the restricted $n$-shot sampling (i.e. $n$ instances instead of $n$ images for each categories) as the novel categories; (ii) while in works such as Meta-DETR [51], the base categories are not few-shot. Both strategies are acceptable for the FSOD task; we adapt to the later one. As shown in  Tab. 2, our method outperforms previous works.
@@ 4.5 Ablation Study
Rich text length. We first verify that the length of the rich text has a distinct impact on the model’s performance. We conduct experiments on the UODD dataset using 4 cases of text: (i) the raw category name, (ii) our manually-built rich text referencing Wikipedia, (iii) an extended version of (ii) by adding image context, (iv) LLM-generated rich text. Cases (ii) to (iv) all result in long sentences that have rich semantics, and thus are generally beneficial for the model performance. See  Tab. 3 for our examples and quantitative results. However, not all extensions on the sentence length are effective. For example, in case (iii), our extended rich text for sea cucumbers has a slightly negative impact over the non-extended version, decreasing performance from 17.4 to 16.9. This may be because the extra image context with an distinct name of another category in UODD (i.e. sea urchin) has confused the model.
Module analysis. In addition, we analyze the effectiveness of our proposed multi-modal aggregated feature module and the rich semantic rectify module. As seen in  Tab. 4, comparing to the single-modal baseline, our method can achieve significant improvement through using rich text multi-modal information. Our rich text rectify module also leads to an additional improvement.
Which modality dominates the result. We conducted this ablation study in  Tab. 5. When only the image modality is utilized, the model corresponds to the baseline configuration. When only the language modality is incorporated, we leverage the CLIP language branch to extract prototype features. Our findings indicate that relying solely on the language modality to generate support prototypes leads to a performance decrease. The reason is that the language modality is prepared for each category, instead of for each image, and thus cannot provide a task-level support feature that is typically seen in general meta-learning settings. However, in  Tab. 4, with the help of the language prototype, we achieve great improvement, as CLIP provides strong vision-language modality feature alignment.
Shared vs decoupled attention. In our MM Aggre. module, we use shared attention module. Alternatively, we could use decoupled attention module, hence, we compare their performance. For decoupled attention, we separate the multi-head self-attention module for the support vision and semantic prototypes. As shown in  Tab. 6, the model’s performance deteriorated with decoupled attention modules. As mentioned in  Sec. 3.2, the language prototypes are designed at the category level, not the image level. Thus, for the task-level support set, a separate attention module tends to diminish the gain from the language modality.
@@
Table 4: The effect of different modules.
Method ArTaxOr DIOR UODD
Image modality only 14.5 19.9 9.2
Language modality only 11.3 7.7 5.1
Table 5: Ablation on single modules.
Method ArTaxOr DIOR UODD
Shared attention 61.4 31.4 17.5
Decoupled attention 46.1 17.7 12.4
Table 6: Ablation on shared vs decoupled attention.
@@ 4.6 Visualization
In  Fig. 5, we present representative qualitative detection results on the three CD-FSOD benchmarks. In these examples, the objective is to detect araneae (ArTaxOr), sea urchins and sea cucumbers (UODD), and windmills (DIOR). The inference uses a confidence threshold setting of 0.3 for all methods. Compared to baseline methods, we show qualitative improvements in a more accurate bounding box (first row), fewer false positives (second row), and improved detection confidence (third row).
@@ 4.7 Discussion
We observe that it is desirable when the rich text description of the target objects are available, but the specific context of the rich text plays a critical role. Consequently, in practice, the user must carefully evaluate the rich text and substantially tune it for greater improvement. This might become a burden in some cases when (i) the user has a rich text, but it contains many descriptions that are irrelevant to the visual information, or (ii) the user does not have rich text and thus need to build one from scratch. For the first case, a general suggestion for practical usage is to keep the information most related to the appearance of the target objects. For the second case, a user could extend and improve upon our LLM-generated text approach. In future work, for example, we plan to train a discriminator to select the best rich text from multiple candidates.